# A Mess of Memory System Benchmarking, Simulation and Application Profiling

### v1

@@ V-C Application profiling
ProfileMe [26] and PerfMemPlus [27] determine whether the application is memory bound based on its memory access latency, measured with the Intel’s Event Based Sampling (PEBS) [28]. The Roofline model [29, 30] analyzes compute performance and memory bandwidth. The application is classified as compute or memory bound based on the comparison with the performance roofs of the target hardware platform. The Top-down model [31] analyzes the application CPI stack. The application is categorized as memory bound if its significant CPI component is caused by the main memory accesses. The model also distinguishes between the memory latency and bandwidth stalls depending on the occupancy of the memory controller queues. Below 70% occupancy, the stalls are considered latency-bound; above this threshold they are bandwidth-bound.
As a part of future work, we will synthesize the Mess application profiling with the Roofline and Top-down models. The Mess and Roofline integration will connect the application memory bandwidth–latency analysis with its compute performance. The Top-down model will provide the CPI stack. Also, the Top-down distinction between the memory latency and bandwidth stalls will be enhanced with the Mess detailed bandwidth–latency analysis.
VI Conclusions
The Memory stress (Mess) framework offers a comprehensive and unified approach to memory system benchmarking, simulation, and application profiling. It is publicly-released and ready to be used by the community for better understanding of the current and exploration of future memory systems.
The Mess benchmark provides a detailed and holistic and memory systems performance view represented by a family of bandwidth–latency curves. The current Mess benchmark release covers x86, ARM, Power, RISC-V, and NVIDIA’s PTX ISAs, and we used it to characterize servers from Intel, AMD, IBM, Fujitsu, Amazon, and NVIDIA with DDR4, DDR5, HBM2, and HBM2E main memory. The Mess characterization significantly expands coverage of the existing tools, uncovering new insights into the memory systems performance and behavior. We also used Mess to benchmark established publicly-available memory models and simulators. Our detailed evaluation shows that the existing memory models poorly resemble the actual system performance, leading to high error rates when simulating memory intensive workloads.
The Mess memory simulator couples memory simulation to memory performance benchmarking and predicts the memory access latencies based on the application position on the bandwidth–latency curves. A memory system can be, therefore, simulated as soon as its bandwidth–latency curves are available, which removes the current lag between the emergence of memory technologies and development of the reliable simulation models. The Mess simulator closely matches the actual memory systems performance and it shows an unprecedented low error of 0.4–6% when simulating widely-used memory benchmarks. It is also fast and reduces the simulation time of cycle-accurate memory simulators by 13–15$\times$. The Mess simulator is easily integrated with ZSim, gem5, and OpenPiton Metro-MPI CPU simulators and it supports a wide range of memory systems based on DDR4, DDR5, Optane, HBM2, HBM2E and CXL memory expanders.
The Mess application profiling determines application position on the corresponding memory bandwidth–latency curves. This information, correlated with other runtime activities and source code, enhances overall understanding of application behavior. The integration of Mess application profiling into a suite of production HPC performance analysis tools further enhances its utility and accessibility.

### v4

@@ VII-C Application profiling
ProfileMe [26] and PerfMemPlus [27] determine whether the application is memory bound based on its memory access latency, measured with the Intel’s Event Based Sampling (PEBS) [28]. The Roofline model [29, 30] analyzes compute performance and memory bandwidth. The application is classified as compute or memory bound based on the comparison with the performance roofs of the target hardware platform. The Top-down model [31] analyzes the application CPI stack. The application is categorized as memory bound if its significant CPI component is caused by the main memory accesses. The model also distinguishes between the memory latency and bandwidth stalls depending on the occupancy of the memory controller queues.
VIII Conclusions
The Memory stress (Mess) framework offers a comprehensive and unified approach to memory system benchmarking, simulation, and application profiling. The Mess benchmark provides a detailed and holistic view of memory systems performance. It covers x86, ARM, Power, RISC-V, and NVIDIA’s PTX ISAs, and it is already deployed to characterize servers from Intel, AMD, IBM, Fujitsu, Amazon, and NVIDIA with DDR4, DDR5, HBM2, and HBM2E main memory. The Mess analytical memory simulator is integrated with ZSim, gem5, and OpenPiton Metro-MPI CPU simulators and it supports a wide range of memory systems, including the CXL memory expanders. It closely matches the actual memory systems performance and it shows an unprecedented low error of 0.4–6% when simulating widely-used memory benchmarks. Finally, the Mess application profiling is already integrated into a suite of production HPC performance analysis tools, adding a new layer of information related to application’s memory performance metrics. The Mess framework is publicly-released and ready to be used by the community for better understanding of the current and exploration of future memory systems.
IX Acknowledgment
This work was funded by the Collaboration Agreement between Micron Technology, Inc. and BSC. The results have been partially funded by the Spanish Ministry of Science and Innovation MCIN/AEI/10.13039/501100011033 (contracts PID2019-107255GB-C21, PCI2022-132935, PCI2021-121958, PID2023-146511NB-I00 and CEX2021-001148-S), the Generalitat of Catalunya (contract 2021-SGR-00763, 2021-SGR-00807 and 2021-SGR-01264), European Union (Grant Agreement No. 955606), European Processor Initiative (EPI) with grant agreement No. 800928, 826647 and 101036168, and by the Barcelona Zettascale Lab (BZL) project with reference REGAGE22e00058408992 and co-financed by the Ministry for Digital Transformation and Public Services, within the framework of the Resilience and Recovery Fund – and the European Union – NextGenerationEU. The work is also supported by the Arm-BSC Center of Excellence and the European HiPEAC Network of Excellence. A. Armejach is a Serra Hunter Fellow. We would also like to acknowledge valuable comments from the following individuals: Julian Pavon (BSC), Mohammad Bakhshalipour (CMU), Vicenc Beltran (BSC), Arash Yadegari (BSC), Nam Ho (FJZ) and Carlos Falquez (FJZ), Stepan Vanecek (TUM), John McCalpin (TACC).
-A Mess benchmark implementation
In this appendix, we discuss Mess benchmark implementation. It includes implementation of pointer-chase and traffic generator. The scripts and manuals for the experimental setup, execution and measurements is included in the public Mess repository. The measurements post-processing removes the outliers, mitigates the noise and plots the results.
-A1 Pointer-chase
The pointer-chase contains a sequence of dependent back-to-back load instructions that access the main memory. Since the instructions are dependent, their execution is serialized, so the average memory access latency is calculated as the ratio between the total execution time and the number of instructions executed. This is guaranteed by the pointer-chase design. In the benchmark initialization, we allocate a contiguous section of memory and initialize it in such a way that a given array element contains the address of the next array element (memory location) that we want to access. This ensures the dependency between the consecutive load instructions. To force accesses to the main memory (and not on-chip caches), the pointer-chase traverses the whole array whose size exceeds the last-level cache. To avoid any cache-line spatial locality, each array element occupies the whole cache line (64 Byte). Finally, to diminish the impact of data prefetching and temporal locality during traversing, the pointer-chase traverses the array in a random pattern. The pointer-chase source code is detailed in Listing 1:
  • Lines 0001–0004 initialize the benchmark: set the number of load instructions, initialize the registers, etc.
  • Lines 0006–1007 are the core of the benchmark. They are a series of back-to-back load instructions for traversing the array. Each load instruction has data dependency on the previous one, so their execution is serialized.
  • Line 1008 decrements the loop counter.
  • Line 1009 checks the loop exit condition.
Listing 1: Pointer-chase source code. x86 assembly. In our study we also use the benchmark for the ARM, IBM Power, and RISC-V ISA.
Line Source code Explanation
0001 register struct line *next asm("rax"); struct line owns pointer to the next access
0002 register int i asm("ecx"); ecx is the loop counter
0003 i = 1000000; initialization of the loop counter. C format.
0004 next = ptr->next first memory access in C format
0005 start_loop: beginning of the loop
0006 mov (%rax), %rax load instruction (pointer-chase)
0007 mov (%rax), %rax load instruction (pointer-chase)
1007 mov (%rax), %rax load instruction (pointer-chase)
1008 dec %ecx decrement loop counter
1009 jnz start_loop if counter is not zero jump to start of the loop
-A2 Traffic generator
We write the core of the memory traffic generator directly in assembly. This enables the development of more precise code (changes at the level of the assembly instruction) and prevents any system software optimizations (e.g. compiler optimizations). The benchmark generates different read and write memory traffic by executing a different mix of load and store instructions in the source code, each of them in a separate assembly file. In the current implementation, the benchmark can execute from 100% load to 100% store operations, with a step of 2%. [^7] The ratio of loads and stores executed by the benchmark does not lead to the same ratio of the reads and writes in the memory traffic. This is because most of the state-of-the-art HPC servers deploy a write-allocate cache policy. With this policy each store operation targeting data in main memory is first read into the cache (memory read), then modified, and finally written to the main memory once the cache line is evicted. Since each store operation causes one memory read and one memory write, e.g. 100% store kernel causes 50%-read 50%-write memory traffic. The generated memory bandwidth is adjusted the issue rate of the memory operations. This is done by interleaving the load/store operation sequence with a configurable dummy loop of nop operations. To reduce the overhead of the loop management, loop counter decrement and jump operations, we employ a long loop of around 100 load/store operations. Listings 2, and 3 show the memory traffic generator implementation in the x86 ISA. The benchmark is also available for ARM, IBM Power, and RISC-V ISAs.
  • Lines 001–008 initialize the benchmark:
    • Lines 001–002 loads the addresses of the arrays a and c into the registers.
    • Lines 003–004 initialize a register as the loop counter.
    • Lines 005–006 set the size of the array a and c.
    • Lines 007–008 set the iteration of the dummy nop function. This controls the rate of the memory operations, i.e. used memory bandwidth.
  • The main loop of the kernel starts at line 009. The loop comprises a series of load and store instructions (Lines 010–015) followed by the call to the dummy nop function (Lines 016–017).
    • The nop function is shown in Listing 3. The number of nop operations is configurable: the larger the number of nop operations in each function call, the lower the rate of memory instruction, and therefore the lower used memory bandwidth. This configuration is managed by the value of nopCount parameter. When nopCount parameter is set to 0, i=0 in Listing 3, the function is finalized without entering the nop loop. This causes only a negligible halt in the load and store sequence, leading to the maximum pressure to the memory system. As we increase the value of the nopCount parameter, we increase the number of the iterations in the dummy nop loop, and therefore reduce the rate of the memory operations issued by the memory traffic generator.
  • The sequence of interleaved load and store instructions, and calls to the nop function is repeated until the line 143. This is the core of the benchmark. In the current implementation, each iteration of the main loop comprises 100 load and store instructions.
  • Lines 144–146 finalize the loop iteration: increment the loop counter, check the exit condition, and conditionally jump to the beginning of the loop.
Listing 2: Memory traffic generator, x86 ISA. Different read and write memory traffic is generated by executing a different mix of load and store instructions in the source code, each of them in a separate assembly file. The generated memory bandwidth is adjusted the issue rate of the memory operations. This is done by interleaving the load/store operation sequence with a configurable dummy loop of nop operations.
Line Source code Explanation
001 register double *a asm("r10"); load array a address into the register r10
002 register double *c asm("r11"); load array c address into the register r11
003 register ssize_t i asm("rbx"); register rbx holds the loop counter
004 i = 0; initialize the counter (i) by 0
005 register ssize_t n asm("r12"); r12 holds the size of the arrays
006 n = *array_size; load r12 with size of the arrays
007 register int *p asm("r15"); register r15 holds the nopCount value
008 p = nopCount; load r15 with value of nopCount
009 ..L_50: the beginning of the kernel loop
010 vmovupd (%r10,%rbx,8), %ymm0; load instruction
011 vmovupd %ymm1, (%r11,%rbx,8); store instruction
012 vmovupd 32(%r10,%rbx,8), %ymm0; load instruction
013 vmovupd %ymm1, 32(%r11,%rbx,8); store instruction
014 vmovupd 64(%r10,%rbx,8), %ymm0; load instruction
015 vmovupd %ymm1, 64(%r11,%rbx,8); store instruction
016 movq %r15, %rdi; send number of nopCount as an input of nop function
017 call nop; call nop function
138 vmovupd 1536(%r10,%rbx,8), %ymm0; load instruction
139 vmovupd %ymm1, 1536(%r11,%rbx,8); store instruction
140 vmovupd 1568(%r10,%rbx,8), %ymm0; load instruction
141 vmovupd %ymm1, 1568(%r11,%rbx,8); store instruction
142 movq %r15, %rdi; send number of nopCount as an input of nop function
143 call nop; call nop function
144 add \$200, %rbx; increment the loop counter
145 cmp %r12, %rbx; check the exit condition
146 jb ..L_50; conditional jump to beginning of the kernel loop
Listing 3: Dummy nop function is used to generate a pause in the sequence of the load and store instructions in the Memory traffic generator (see Listing 2). The input parameter nopCount is the number of the dummy loop iterations. The larger the number of iterations, the lower the rate of memory instruction and the lower the generated memory bandwidth.
Line Source code Explanation
001 int nop(int *nopCount) { definition of nop function
002 unsigned int i = *nopCount; i is the loop counter; number of loops for nop operations
003 if ( !i ) { if i is zero then return immediately
004       return 0;
005 } else { assembly section of nop function
006       asm( it is written in extended assembly format
007          mov %0, %%ecx; move the value of %0 register into ecx register
008          the_loop%=: the beginning of the loop
009          nop;        nop instruction
010          dec %%ecx        decrementing the counter
011          jnz the_loop%=        conditional jump to beginning of the loop
012          : "r" (i) : ecx put the value of i into register %0 (extended assembly)
013       );
014 }
014 return 0;
015 }
-B Case study: Remote-socket memory emulation of CXL memory expanders
Recent industrial studies emulate CXL memory expansion in cloud and datacenter servers with conventional dual-socket systems: one socket is used as a host CPU and the other socket as a CPU-less memory expander [102, 103]. To evaluate this approach, we perform a detailed simulation of both systems. The CPU host is simulated with the ZSim configured to model the Intel 24-core Skylake processor (Section IV-B). We use Micron Technology’s bandwidth–latency curves to model the CXL memory expander, and measurements from the actual dual-socket server to model the remote-socket memory. In each system configuration, we simulate 500 billion instructions of the multiprogramed SPEC CPU2006 workloads [104].
Figure 17 plots the CXL (green) and remote-socket (blue) memory curves, and the benchmark behavior. Each 20.000 simulated instructions we monitor the benchmark read and write memory bandwidth, and then plot each observation as a point on the corresponding bandwidth–latency curve. The figure shows two characteristic use-cases. The perlbench benchmark (Figure 17(a)) has a low bandwidth utilization. In this area remote-socket memory curves shows approximately 28 ns higher latency, which leads to somewhat lower performance w.r.t. target CXL system. We detect the opposite behavior for high-bandwidth benchmarks. The remote-socket curves have a higher bandwidth saturation area. This leads to higher achieved bandwidths and the overall performance of the bandwidth-intensive benchmarks, such as the lbm illustrated in Figure 17(b). The overall performance trend is illustrated in Figure 18. The figure plots the results for all SPEC CPU2006 benchmarks, sorted from the lowest to the highest bandwidth utilization. For low-bandwidth benchmarks, remote-socket emulation provides up to 12% lower performance w.r.t. to the target CXL system. Both system provide a similar performance for the benchmarks with medium bandwidth utilization between 30% and 50% of CXL max theoretical bandwidth. For the high-bandwidth benchmarks, the remote-socket memory provides 11%–22% higher performance.
To the best of our knowledge, this is the first study that enables a detailed simulation of CXL memory expanders based on the manufacturer’s SystemC model. The model is already integrated and publicly released with ZSim, gem5 and OpenPiton Metro-MPI simulators. We also provide the performance analysis and correction factors that can be useful for future studies who want to keep useing remote-socket memory emulation, e.g. due to its much faster execution time.
(a) The perlbench experiences higher remote-memory latency which leads to 5% lower performance.
(b) The lbm reaches higher remote-memory bandwidth which leads to 11% higher performance.
Figure 17: Remote-socket emulation of CXL memory expanders: Bandwidth–latency curves and implications
Figure 18: Remote-socket memory emulation of CXL memory expanders: Application performance difference is correlated with its bandwidth utilization. SPEC2006 benchmarks, 24-core Intel Skylake CPU.
-C Artifact Appendix
-C1 Abstract
This artifact comprises the implementation of the Mess benchmark for various platforms, including actual hardware platforms (e.g., Intel, IBM, NVIDIA), system simulators (ZSim, gem5, and OpenPiton), and memory simulators (DRAMsim3, Ramulator, Ramulator2). It also includes the Mess memory simulator integrated with ZSim, gem5, and OpenPiton simulators.
This artifact also contains all the scripts and guidelines necessary to reproduce the major figures presented in this paper. In addition to the scripts, it includes raw hardware measurements, processed measurements, and the final bandwidth-latency curves.
This study involves more than eight different hardware platforms with a diverse set of runtime environments, including various compilers, ISAs, and tools. Therefore, in this artifact, we mention the packages, tools, and applications without specifying the exact versions. In the Git repository, we provide detailed dependencies and version information for each experiment.
-C2 Artifact check-list (meta-information)
  • Program: Pointer-chase and workload generator implemented in C, C++ with its kernel implemented in inline assembly (included in the benchmark). Mess simulator integrated with ZSim, gem5, and OpenPiton.
  • Compilation: GCC, G++, ICX, MPI++, and Python 3.
  • Data set: All the raw values measured from hardware counters and simulation tracecs as well as final curves for all the figures are included.
  • Run-time environment: For Intel Cascade Lake curves, we use small server with single node. For Fujitsu A64FX, we use PJM batch processing support. For Graviton server, we use Amazon AWS. For the rest of the systems, we use production servers with Slurm Workload Manager environment.
  • Hardware: Servers or Supercomputers with the following CPUs and GPUs: Intel Sandy Bridge, Skylake, Cascade Lake and Sapphire Rapids. AMD Zen 2, IBM Power 9, Amazon Graviton 3, Fujitsu A64FX, NVIDIA Hopper H100.
  • Metrics: For latency, we use nanoseconds and cycles. For bandwidth, we use GB/s.
  • Output: We plot bandwidth–latency curves. Moreover, we print detailed data points in a .csv file.
  • Experiments: Generate experiments using supplied scripts.
  • How much disk space required (approximately)?: 10s of GB.
  • How much time is needed to prepare workflow (approximately)?: For each experiment, approximately one hour.
  • How much time is needed to complete experiments (approximately)?: To generate bandwidth–latency curves for actual hardware, ZSim, gem5, and OpenPiton, we need approximately 3-6 days, 1-2 weeks, 2-3 weeks, and 1-2 weeks, respectively.
  • Publicly available: Yes.
  • Code licenses: MIT License.
  • Data licenses: MIT License.
  • Archived (provide DOI)?: 10.5281/zenodo.13748673
-C3 Description
Figure 19 and 20 show where to replicate each major result presented in the paper. A detailed explanation of dependencies, system configurations, experimental setups, and result validation is available in the README.md file in each folder. To fit within the page limit of this artifact, this guideline introduces the general approach for replicating each result, along with an example to reproduce Figure 6.b. To replicate other figures, a similar approach should be followed (detailed guidelines are available in the Git repositories).
How to access
The Mess Benchmark artifact can be cloned from GitHub at https://github.com/bsc-mem/Mess-benchmark.git. The structure of the repository is depicted in Figure 19. The Mess simulator artifact can also be cloned from https://github.com/bsc-mem/Mess-simulator. The structure of the repository is depicted in Figure 20. Each folder in the repositories replicates one or more figures presented in the main manuscript (figures are indicated in blue text). This artifact can also be downloaded as a .zip file from https://zenodo.org/records/13748674.
Hardware dependencies
To run the Mess benchmark on actual hardware, access to a full node is required. The CPU/GPUs must support hardware counters to measure memory bandwidth (preferably uncore counters). For simulation experiments, a single core is sufficient. However, ZSim and OpenPiton can benefit from multicore or multinode parallelism.
Software dependencies
The benchmark and simulations run on Linux OS. To measure uncore counters, we primarily use the Linux perf tool, which is supported by all major Linux versions. In some cases, we also use Intel VTune and LIKWID.
Data sets
All the data sets are included in the repositories.
-C4 Installation
To install, first clone the repository. Then, navigate to the directory corresponding to the figure from the main manuscript that you want to replicate (see Figures 19 and 20). For Figure 6.b of the main manuscript:
git clone https://github.com/bsc-mem/Mess-benchmark.git
cd ./CPU/Simulators/Trace-driven/DRAMsim3
[…]
.1 Mess-benchmark. .2 Actual-hardware. .3 CPU. .4 x86 (Fig.2.a,b,c,f). .4 ARM (Fig.2.e,g). .4 IBM power (Fig.2.d). .4 RISC-V. .3 GPU. .4 NVIDIA PTX (Fig.2.h). .2 Simulators. .3 Execution-driven. .4 gem5 (Fig.4.b,c,d). .4 ZSim (Fig.5.b–f). .3 Trace-driven. .4 Ramulator2 (Fig.6.a). .4 DRAMsim3 (Fig.6.b). .4 Ramulator (Fig.6.c).
Figure 19: Mess benchmark repo
[…]
.1 Mess-simulator. .2 Standalone. .2 Integrated. .3 ZSim
(Fig.10.a–c)
. .3 gem5
(Fig.11.a–b)
. .3 OpenPiton
(Fig.14.b).
Figure 20: Mess simulator repo
[…]
Figure 21: Directory structure to replicate DRAMsim3 experiments.
-C5 Experiment workflow
For each experiment (each folder in Figures 19 or 20), the workflow for running the Mess benchmark/simulation is provided in the runner.sh script. For Figure 6.b of the main manuscript (trace-driven DRAMsim3 simulation), the full workflow takes less than one day.
-C6 Evaluation and expected results
To replicate each experiment, we run replicate.sh script. This script compiles necessary codes, executes the workflow (i.e., runner.sh), and processes the output raw data. The artifact also includes all the raw measurements and final processed .csv data. For the DRAMsim3 example, the replicate.sh script inside DRAMsim3 directory (Figure 21) executes the following commands:
# unzip trace files
cd traceInput
for file in *.zip; do
unzip ”\$file”
done
cd ..
# compile DRAMsim3
cd DRAMsim3_mn5
mkdir build
cd build
cmake ..
make
cd ../..
# run the experiment
./runner.sh
# Post-processing #
# generate results.csv file
python3 main.py .
# generate output.pdf (bandwidth–latency curves)
python3 convert.py
The easiest way to validate the result is visually by examining the generated curves (e.g., output.pdf in our example). However, if one wants to evaluate the results in more detail, the results.csv file can be compared to results_original.csv; rows with the same rw_ratio and pause values should have a very close latency and bandwidth.